5.30 Suppose two random variables Xand Yare both zero mean and unit variance. Furthermore, assume they have a correlation coefficient of p. Two new random variables are formed according to W = aX + bY, Z = cX + dY. Determine under what conditions on the constants a,b. c, and d the random variables W and Z are uncorrelated.
Answer #1

Here

E(X) = E(Y) = 0

and Var(X) = Var(Y) = 1

and corr(X,Y) = \rho \Rightarrow cov(X,Y) = \rho

Cov(W,Z) = Cov(aX+bY, cX+dY)

\Rightarrow Cov(W,Z) = acVar(X)+adCov(X,Y)+bcCov(X,Y)+bdVar(Y)

\Rightarrow Cov(W,Z) = ac*1+ad*\rho +bc*\rho +bd*1

\Rightarrow Cov(W,Z) = ac+bd+\rho (ad+bc)

Now,

Cov(W,Z) = 0

\Rightarrow ac+bd+\rho (ad+bc)= 0

\Rightarrow \rho (ad+bc)= -(ac+bd)

\Rightarrow \rho= \frac{-(ac+bd)}{(ad+bc)}

If a,b,c and d are such that \rho= \frac{-(ac+bd)}{(ad+bc)} ,then W and Z will be uncorrelated.

Moreover, ac = -bd and ad = -bc,then also W and Z will be uncorrelated.